Solve the system by addition
–4x + 5y = 1
x – 4y = 19

Answer by sdmmadam@yahoo.com(530)

Solve the system by addition
–4x + 5y = 1 ----(1)
x – 4y = 19 ----(2)X4
The stipulation is to solve by addition
Observe that one coefficient of x is (-4) and the other is 1
So multiply equation (2) by 4
4x-16y= 76 ----(3)
And –4x + 5y = 1 ----(1)
(3)+(1) implies
(4x-4x)+(-16y+5y) = (76+1)
0-11y = 77
-11y = 77
y = 77/(-11) (7X11)/(-11) = -7 (cancelling 11)
y = -7
Putting y = -7 in (2)
x-4y =19
x-4X(-7) = 19
x+28 = 19
x = 19-28
x=-9
Answer: x= -9 and y =-7
Verification: Putting x= -9 and y =-7 in the other equation, that is in (1)
–4x + 5y = 1 ----(1)
LHS= -4x+5y =-4X(-9)+5X(-7) = 36-35 = 1 =RHS
Therefore our values are correct.